Find the slope and y-intercept of the equation
y=-1/7x-10

Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

Y=mX+b WHERE m=SLOPE & b=THE Y INTERCEPT.
m=-1/7 IS THE SLOPE.
b=-10 IS THE Y INTERCEPT.

Answer by tutorcecilia(2152)

In the slope-intercept formula: y=mx+b
y=any y-value
m=the slope
x=any corresponding x-value
b=the y-intercept
.
Therefore, the slope m=-1/7 and the y-intercept = -10
y=-1/7x-10